Eva wants to build a fence around her circular garden.if the gardens radius is 12 ft what is the best estimate for the length of the fence ? A. 19 ft B. 24 ft C. 38 ft or D. 75
Answer by stanbon(75887)

Eva wants to build a fence around her circular garden.if the gardens radius is 12 ft what is the best estimate for the length of the fence ?
-----
Perimeter = 2(pi)r
P = 2(pi)*12 = 24(pi) is approximately 75 ft
Cheers,
Stan H.
==============
A. 19 ft B. 24 ft C. 38 ft or D. 75